OpenStudy (anonymous):

A vending machine sells chips at $0.55 and candy at $0.75. Last month, the vending machine yielded $170.00 with the sale of 260 items. let X be the number of chips sold and Y the number of candies. How many packages of chips were purchased last month?

OpenStudy (anonymous):

We know that X + Y = 260 and that 0.55X + 0.75Y = 170

OpenStudy (anonymous):

So X = 260 - Y. now plug it in the other equation and solve.

OpenStudy (anonymous):

or try to multiply first equation with -0.55 and then aggregate with another one